OpenStudy (anonymous):

Find the slope of the tangent line to the ellipse x^2/9+y^2/25=1 at the point (x,y). Are there any points where the slope is not defined? If so, provide those x-values.

OpenStudy (anonymous):

So far I have \[y'=\frac{ -25x }{ 9y }\] I just do not know how to answer the second part of the question

OpenStudy (e.mccormick):

Well, are points of infinite slope or 0 slope defined?

OpenStudy (amistre64):

x^2/9+y^2/25=1 2x/9 + 2y/25 y' = 0 x/9 + y/25 y' = 0 y/25 y' = -x/9 y' = -25x/9y good job on that part :)

OpenStudy (amistre64):

what makes y' not definable?